OpenStudy (shaleiah):

@mdillon1230 I need help with algebra

OpenStudy (anonymous):

i completely suck at algerbra but throw it at me lol

OpenStudy (anonymous):

Let x = one of two real numbers and y - second of two real numbers . Since we have two unknowns, we'll need two equations. . From:"that have a sum of 8" we get equation 1: x + y = 8 . From:"product of 2" we get equation 2: xy = 2 . Solving equation 1 for y: x + y = 8 y = 8-x . Plug the above into equation 2 and solve for x: xy = 2 x(8-x) = 2 8x-x^2 = 2 0 = x^2 - 8x + 2 Since, we can't factor, we must use the quadratic equation. x = {7.742, 0.258} y = (0.258, 7.742) . Answer: the two numbers are 7.742 and 0.258

OpenStudy (anonymous):

how's that.

OpenStudy (shaleiah):

lol i just needed the written expression not the solution and quadratic.

OpenStudy (anonymous):

oh lol x + y = 8 y = 8-x

OpenStudy (anonymous):

feel free to give medals xD

OpenStudy (shaleiah):

its incorrect

OpenStudy (anonymous):

The sum of eight 8 and the product of two 8+2* and a number x. 8+2*x In traditionalform: 2x+8

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

im sorry @shaleiah i said i suck at algebra now we know to tagg @robtobey